OpenStudy (anonymous):

Let me ask this again... Simplify (15/25)^-3 I can only simplify it up to 1/(3/5)^3, which is where I am stuck. Is this as far as it can go, or is there more to it? Please help!

OpenStudy (anonymous):



OpenStudy (anonymous):

^ This is as far as I can go before I get stuck. Is this the answer or not?

OpenStudy (4n1m0s1ty):

Not quite you are almost there, just look at the denominator first (ignore the numerator). What is (3/5)^3?

OpenStudy (anonymous):

That's where I get confused, because I think I might be doing it wrong.

OpenStudy (4n1m0s1ty):

well whats (3/5)*(3/5)*(3/5)?

OpenStudy (4n1m0s1ty):

or alternatively (3*3*3)/(5*5*5)?

OpenStudy (campbell_st):

simplify the brackets 1st \[\frac{15}{25} = \frac{3}{5}\] using index notation it can be written as \[\frac{3}{5} = 3^1 \times5^{-1}\] using the power of a power law \[(x^a)^b = x^{a \times b}\] the problem is \[(3^1 \times 5^-1)^{-3} = 3^{1 \times -3} \times 5^{-1\times -3}\] which gives \[3^{-3} \times5^3\] using fraction notation is \[\frac{1}{3^3} \times 5^3 = \frac{5^3}{3^3} = \frac{125}{27}\]

OpenStudy (4n1m0s1ty):

or just read campbell's answer

OpenStudy (anonymous):

OK, thanks!